Problem 14: An electric dipole consists of a particle with a charge of -6.0 x 106 C at the origin and a particle with a chargPlease explain problems 14 & 15.
Answer #1

14

P = qD = 6* 10^-6* 3* 10^-3 = 1.8* 10^-8 Cm

directed along - y axis

option (e) is correct

======

15. option (c) is correct

because two diametrically opposite charges contributes zero field, hence only field is given by bottom left charge towards it

=====

Comment before rate in case any doubt, will reply for sure.. goodluck